### v1

# Optimal design of unimorph-type cantilevered piezoelectric energy harvesters using level set-based topology optimization by considering manufacturability

Abstract
In this study, we proposed a design methodology for a piezoelectric energy-harvesting device optimized for maximal power generation at a designated frequency using topology optimization. The proposed methodology emphasizes the design of a unimorph-type piezoelectric energy harvester, wherein a piezoelectric film is affixed to a singular side of a silicon cantilever beam. Both the substrate and the piezoelectric film components underwent concurrent optimization. Constraints were imposed to ensure that the resultant design is amenable to microfabrication, with specific emphasis on the etchability of piezoelectric energy harvesters. Several numerical examples were provided to validate the efficacy of the proposed method. The results showed that the proposed method derives both the substrate and piezoelectric designs that maximize the electromechanical coupling coefficient and allows the eigenfrequency of the device and minimum output voltage to be set to the desired values. Furthermore, the proposed method can provide solutions that satisfy the cross-sectional shape, substrate-depend, and minimum output voltage constraints. The solutions obtained by the proposed method are manufacturable in the field of microfabrication.
keywords:
Topology optimization, Piezoelectric material, Energy harvester, Manufacturability, Level set method
[…]
[inst0]organization=Osaka Research Institute of Industrial Science and Technology,addressline=Ayumino–2–7–1, city=Izumi–city, postcode=594-1157, state=Osaka, country=Japan
[…]
[inst1]organization=Department of Mechanical Engineering, Graduate School of Engineering, The University of Tokyo,addressline=Yayoi 2–11–16, city=Bunkyo–ku, postcode=113–8656, state=Tokyo, country=Japan
[…]
[inst2]organization=Department of Strategic Studies, Institute of Engineering Innovation, Graduate School of Engineering, The University of Tokyo,addressline=Yayoi 2–11–16, city=Bunkyo–ku, postcode=113–8656, state=Tokyo, country=Japan
1 Introduction
Piezoelectric energy harvesters convert mechanical energy into electrical energy through the piezoelectric effect. These devices have gained significant attention in recent years owing to their potential to scavenge energy from various sources such as vibrations Jeon et al. (2005); Roundy et al. (2003) and fluid flow Yun and Kim (2016). The design of a piezoelectric energy harvester plays a critical role in determining its efficiency and performance Erturk and Inman (2011). Various piezoelectric energy harvester designs, including cantilever beams, multilayered structures, and cylinder structures, have been proposed and studied Cook-Chennault et al. (2008); Beeby et al. (2006). The unimorph cantilevered piezoelectric energy harvester has several advantages over the other piezoelectric energy harvester designs. The simple and compact design of this harvester makes it suitable for applications where size and weight are critical factors, such as in portable or wearable devices. Furthermore, the unimorph cantilevered design allows easy integration with other systems such as sensors or wireless transmitters Saxena et al. (2021). However, the performance of a unimorph cantilevered piezoelectric energy harvester is limited by the amount of mechanical energy harvested from external vibrations. To overcome these limitations, researchers have explored the use of advanced materials and optimization techniques to improve the performance of unimorph cantilevered piezoelectric energy harvesters. Some studies have investigated piezoelectric materials with improved electromechanical coupling factors to enhance the piezoelectric effect and improve energy conversion efficiency Yoshimura et al. (2013); Murakami et al. (2013). Other researchers have focused on designing the geometry of the cantilever beam and piezoelectric element for maximum energy output. For instance, rectangular, triangular, and trapezoidal shapes have been proposed for unimorph piezoelectric vibration energy harvesters Glynne-Jones et al. (2001). However, as these structures have simple geometries, significant improvements can be achieved using topological optimization techniques Bendsøe and Kikuchi (1988).
[…]
Sato et al. Sato et al. (2017), Yamada et al. Yamada and Noguchi (2022), and Tajima et al. Tajima and Yamada (2023) developed advanced methodologies that leverage fictitious physical models to assess manufacturability. Sato et al. Sato et al. (2017) developed a method for evaluating manufacturability in molding processes using fictitious physical models to ensure that the geometrical features of mold parts are amenable to manufacturing. Yamada et al. Yamada and Noguchi (2022) focused on additive manufacturing and proposed a topology optimization method that addressed closed cavity exclusion constraints and enhanced the design process of additive manufacturing. Tajima et al. Tajima and Yamada (2023) introduced a coupled fictitious physics model aimed at improving the convergence in topology optimization by integrating the objectives of both the fictitious physical and mechanical models, particularly for additive manufacturing with geometric constraints. These methodologies collectively underscore the importance of incorporating manufacturability considerations into topology optimization, which is essential for practical applications in various manufacturing processes. The level set-based topology optimization championed by Yamada et al. Yamada et al. (2010) affords precision in modulating the geometrical intricacy of the resultant optimal configurations. Building on this foundation, Noda et al. Noda et al. (2022) augmented this level set strategy to cater to the nuances of multi-material topology optimization. This innovation ensures a meticulous orchestration of the complexity intrinsic to each deployed material, culminating in a multi-material assembly that is conducive to straightforward manufacturing. However, the previously underscored methodologies are predominantly conceptualized for traditional, macroscopic machining techniques such as milling. Therefore, their direct application to piezoelectric energy harvesters, typically realized through microfabrication techniques such as etching Murakami et al. (2013); Aramaki et al. (2019), may not be feasible.
In this study, we introduced a methodology for the optimal design of piezoelectric energy harvesters by employing topology optimization anchored in the level set method. This approach emphasizes manufacturability in the field of microfabrication, with etching being the predominant process under consideration. The design criteria for piezoelectric energy harvesters were delineated based on the operational frequency and a threshold for the excitation voltage output. Furthermore, the proposed method uniquely facilitates the concurrent optimization of three-dimensional structures within both the substrate and piezoelectric material domains. The resulting designs not only satisfy the stipulated specifications but are also readily adaptable to microfabrication procedures.
The remainder of this paper is organized as follows: In Section 2, the architectural framework of the designated device is described. After this initial exposition, in this section we further elaborate on the optimal design approach, articulating the pivotal design criteria, including the eigenfrequency specifications and constraints regarding the minimum output voltage. In Section 3, we explain the formulation of constraints pertinent to manufacturability. Specifically, we introduced two constraints to enhance manufacturability: ensuring a consistent cross-sectional shape across each domain and the stipulation that no piezoelectric material is positioned without an accompanying substrate. In Section 4, implementation of the numerical analysis of the proposed method is described. In Section 5, numerical illustrations by employing a benchmark model to test the efficacy of the introduced method are detailed. Finally, in Section 6, the conclusions of this study are summarized. Additionally, the methodologies tailored for computing the output voltage are detailed in A, and the sensitivity analyses are discussed in B.
2 Formulation for the optimal design of a piezoelectric energy harvester
2.1 Unimorph cantilevered piezoelectric energy harvester
This study focused predominantly on the unimorph cantilevered energy harvester. An unimorph cantilevered piezoelectric energy harvester comprises three main components: substrate, piezoelectric material, and weight. These elements are depicted in Figure 1, which shows the theoretical structure of an unimorph cantilevered piezoelectric vibration energy harvester. The unique advantage of the unimorph design is its capability to produce a superior electrical output for a given mechanical input compared with a conventional cantilevered piezoelectric energy harvester. This superiority is attributed to the mechanical amplification effect induced by the nonpiezoelectric layer, which result in an enhanced deformation of the piezoelectric layer for a specific input vibration.
Figure 1: Concept of a piezoelectric energy harvester
2.2 Formulation of piezoelectric phenomena
The governing equations for a piezoelectric energy harvester are articulated in the following set of simultaneous piezoelectric equations, as elaborated in Erturk and Inman (2011):
$\begin{split}\bm{\sigma}&=\bm{C}^{E}\bm{s}(\bm{u})+\bm{e}^{T}\bm{\nabla}\varphi\\ \bm{d}&=\bm{e}\bm{s}(\bm{u})+\bm{\varepsilon}^{S}\bm{\nabla}\varphi,\end{split}$ (1)
$\bm{s}(\bm{u})=\frac{1}{2}(\bm{\nabla}\bm{u}+(\bm{\nabla}\bm{u})^{T}).$ (2)
Here, $\bm{\sigma}$ is the stress tensor; $\bm{d}$ is the electric displacement tensor; $\bm{u}$ is the displacement vector; $\bm{s}(\bm{u})$ represents the strain tensor; $\bm{C}^{E}$ is the elasticity tensor at a constant electric field; and we assume that $\bm{C}^{E}$ represents an isotropic linear elastic material; $\bm{e}$ represents the piezoelectric stress constant tensor; $\varphi$ is the electric potential; $\bm{\varepsilon}^{S}$ is the dielectric constant tensor at a constant strain; and ${}^{T}$ denotes the transpose of the tensor.
In this study, we primarily focused on a piezoelectric energy harvester under harmonic excitation at a frequency $\omega$. Within this context, the displacement vector $\bm{u}$ and the electric potential $\varphi$ are designated as state variables, whereas $\bm{v}$ and $v_{\varphi}$ serve as test functions. Consequently, Eq. (1) is transcribed into its weak form as follows:
$\begin{split}\int_{\Omega}\bm{s}(\bm{v})^{T}\bm{C}^{E}\bm{s}(\bm{u})\ \text{d}\Omega-\omega^{2}\rho\int_{\Omega}\bm{v}^{T}\bm{u}\ \text{d}\Omega+\int_{\Omega}\bm{s}(\bm{v})^{T}\bm{e}^{T}\bm{\nabla}\varphi\ \text{d}\Omega=\int_{\Omega}\bm{v}^{T}\hat{\bm{f}}\ \text{d}\Omega\\ \int_{\Omega}(\bm{\nabla}v_{\varphi})^{T}\bm{e}\bm{s}(\bm{u})\ \text{d}\Omega-\int_{\Omega}(\bm{\nabla}v_{\varphi})^{T}\bm{\varepsilon}^{S}\bm{\nabla}\varphi\ \text{d}\Omega=\int_{\Omega}v_{\varphi}\hat{\bm{q}}\ \text{d}\Omega.\end{split}$ (3)
Here, $\rho$ refers to the material’s density, $\hat{\bm{f}}$ is the external forcing function, and $\hat{\bm{q}}$ represents the externally provided electric charge. In this study, we assumed steady-state vibration.
The electromechanical coupling coefficient was employed as a metric to assess the performance of the piezoelectric energy harvester. The coefficient is calculated using the following equation, which considers the $i$-th mode eigenfrequency $\omega_{oci}$ under electrode open-circuit conditions and the $i$-th mode eigenfrequency $\omega_{sci}$ under electrode short-circuit conditions Hagood and von Flotow (1991):
$k^{2}_{i}=\frac{\omega_{oci}^{2}-\omega_{sci}^{2}}{\omega_{oci}^{2}}.$ (4)
Under open-circuit conditions (where $\hat{\bm{q}}=0$), the piezoelectric simultaneous equations are expressed as follows:
$\begin{split}\int_{\Omega}\bm{s}(\bm{v}_{oci})^{T}\bm{C}^{E}\bm{s}(\bm{u}_{oci})\ \text{d}\Omega-\omega_{oci}^{2}\rho\int_{\Omega}\bm{v}_{oci}^{T}\bm{u}_{oci}\ \text{d}\Omega+\int_{\Omega}\bm{s}(\bm{v}_{oci})^{T}\bm{e}^{T}\bm{\nabla}\varphi_{oci}\ \text{d}\Omega&=0\\ \int_{\Omega}\bm{\nabla}v_{\varphi_{oci}}^{T}\bm{e}\bm{s}(\bm{u}_{oci})\ \text{d}\Omega-\int_{\Omega}\bm{\nabla}v_{\varphi_{oci}}^{T}\bm{\varepsilon}^{S}\bm{\nabla}\varphi_{oci}\ \text{d}\Omega&=0,\end{split}$ (5)
where $\bm{u}_{oci}$ is the eigenvector corresponding to the eigenvalue $\omega_{oci}$, $\varphi_{oci}$ denotes electric potential under open-circuit condition, and $\bm{v}_{oci}$ is the test function. Conversely, under short-circuit conditions (where $\bm{\nabla}\varphi=0$), the piezoelectric simultaneous equations become solely the equation for harmonic oscillation, unaffected by the influence of the piezoelectric material.
$\int_{\Omega}\bm{s}(\bm{v}_{sci})^{T}\bm{C}^{E}\bm{s}(\bm{u}_{sci})\ \text{d}\Omega-\omega_{sci}^{2}\rho\int_{\Omega}\bm{v}_{sci}^{T}\bm{u}_{sci}\ \text{d}\Omega=0,$ (6)
where $\bm{u}_{sci}$ is the eigenvector corresponding to the eigenvalue $\omega_{sci}$ and $\bm{v}_{sci}$ is the test function.
In the context of piezoelectric energy harvester design, the eigenfrequency of the device is established as a crucial design parameter. Consequently, in this study, we considered strategies for concurrently aligning the eigenfrequency with the desired value. Adopting the objective function proposed by Yamada et al. Yamada et al. (2012), we normalized the eigenfrequency specification problem with respect to the $i$-th mode desired frequency $\bar{\omega}_{i}$ and formulated the objective function $F{\omega}$ as follows:
$F_{\omega}=$ $\sum_{i=1}^{n}\frac{|\omega_{oci}-\bar{\omega}_{i}|^{2}}{\bar{\omega}_{i}^{2}},$ (7)
where we considered $F_{\omega}$ as a problem specifying up to the $n$-th mode.
To transform the maximization problem into a minimization problem, we formulated the objective function $F_{k}$ using the electromechanical coupling coefficient $k_{i}$ as follows:
$F_{k}=\sum_{i=1}^{n}\frac{1}{k^{2}_{i}\bar{\omega}_{i}^{2}}=\sum_{i=1}^{n}\frac{\omega_{oci}^{2}}{(\omega_{oci}^{2}-\omega_{sci}^{2})\bar{\omega}_{i}^{2}}.$ (8)
By weighting the modes by $\bar{\omega}_{i}$, we emphasized the lower-order modes.
In this study, we formulated an optimization problem that simultaneously minimizes distinct objective functions, that is, $F_{sb}$ for the substrate material domain $\Omega_{sb}$ and $F_{pe}$ for the piezoelectric material domain $\Omega_{pe}$, while maintaining a shared state field. This formulation incorporates the weighting factors $\alpha_{sb}$ and $\alpha_{pe}$ as demonstrated below Kim and Lee (2014):
$\begin{split}\underset{\Omega_{sb}}{\text{inf}}\qquad&F_{sb}=\alpha_{sb}F_{k}+(1-\alpha_{sb})F_{\omega}\\ \underset{\Omega_{pe}}{\text{inf}}\qquad&F_{pe}=\alpha_{pe}F_{k}+(1-\alpha_{pe})F_{\omega}\\ \text{subject}\;\text{to}\qquad&\text{governing}\;\text{equations}\;\text{for}\;\text{an}\;\text{open}\;\text{circuit},\\ \qquad&\text{governing}\;\text{equations}\;\text{for}\;\text{a}\;\text{short}\;\text{circuit}.\end{split}$ (9)
2.2.1 Minimum output voltage constraint
Concomitant with the eigenfrequency, the output voltage is a pivotal performance metric in the design of piezoelectric energy harvesters. This attribute is significant because devices harnessing the generated energy often require a minimum viable output voltage. Owing to the physical constitution of the piezoelectric energy harvester, wherein the piezoelectric material is encapsulated between electrodes, the resultant output voltage from the vibratory motion is closely correlated not only with the charge induced on the electrodes but also with the capacitance between the piezoelectric material and the electrodes. A larger piezoelectric material area results in a greater charge, albeit at the cost of increased capacitance. Hence, the capacitance must be meticulously regulated to match the required output voltage. Macroscopically, the output voltage $V_{E}$ is expressed as follows:
$V_{E}=\frac{Q}{C_{p}}=\frac{\int_{\Omega_{pe}}\bm{n}_{z}\cdot\nabla\varphi\ \text{d}\Omega}{\int_{\Omega_{pe}}\ \text{d}\Omega/(\varepsilon_{z}L_{z}^{2})},$ (10)
where $C_{p}$ signifies the capacitance of the piezoelectric energy harvester, $Q$ denotes the electrical charge elicited by the piezoelectric material, $\bm{n}_{z}$ represents a unit vector in the polarization direction, $\varepsilon_{z}$ is the polarization direction component of the piezoelectric material’s dielectric constant tensor, and $L_{z}$ is the thickness of the piezoelectric domain. The constraint on the minimum output voltage is formulated as follows:
$V_{E}=\frac{\int_{\Omega_{pe}}\bm{n}_{z}\cdot\nabla\varphi\ \text{d}\Omega}{\int_{\Omega_{pe}}\ \text{d}\Omega/(\varepsilon_{z}L_{z}^{2})}\geq\bar{V}_{minV},$ (11)
$\int_{\Omega_{pe}}\ \text{d}\Omega\leq\frac{\varepsilon_{z}L_{z}^{2}\int_{\Omega_{pe}}\bm{n}_{z}\cdot\nabla\varphi\ \text{d}\Omega}{\bar{V}_{minV}},$ (12)
where $\bar{V}_{minV}$ denotes the minimum output voltage. This formulation allows the minimum output voltage constraint to be treated similarly to the volume constraint.
2.3 Concept of topology optimization
Topology optimization is a type of structural optimization method. Structural optimization is used to obtain a structure $\Omega$ that minimizes or maximizes an objective function. The objective function often includes physical properties, such as stiffness Bendsøe and Kikuchi (1988); Allaire (2001), as well as thermal Zhou et al. (2019); Yamada et al. (2011); Jing et al. (2015); Wu et al. (2019); Miki and Yamada (2021), fluid Feppon et al. (2020); Guan et al. (2023), electromagnetic Choi and Yoo (2009); Yamada et al. (2013), and acoustic properties Wadbro and Berggren (2006); Sigmund and Jensen (2003); Noguchi and Yamada (2022, 2021). Therefore, the optimal structure is obtained under the assumption that the objective function satisfies the governing equations that describe the physical phenomena. The governing equations are considered as constraints in the optimization problem, and the basic structural optimization problem can be formulated as follows:
$\begin{split}\underset{\Omega}{\text{inf}}\qquad&F(u,\Omega)=\int_{\Omega}f(u)\mathrm{d\Omega}\\ \text{subject to}\qquad&\text{the governing equation system},\\ &\text{constraint equations},\end{split}$ (13)
where $u$ is the state variable obtained as the solution of the governing equation and $f(u)$ is the objective function.
Next, we considered the application of topology optimization to the structural optimization problem (13). We introduced a domain $\Omega_{D}\subset\mathbb{R}^{n}(n=2\;\text{or}\;3)$ where the structure can be placed. Here, the domain $\Omega_{D}$ is called a fixed design domain because it does not change during the optimization process. The fixed design domain consists of a domain filled with the structure (hereafter referred to as the material domain) and a domain not filled with the structure (hereafter referred to as the void domain), and these domains are expressed by the characteristic function $\chi$, which is defined as follows:
$\chi(\bm{x}):=\left\{\begin{array}[]{ll}1\qquad\mathrm{for}\quad\bm{x}\in\Omega\\ 0\qquad\mathrm{for}\quad\bm{x}\in\Omega_{D}\backslash\Omega,\\ \end{array}\right.$ (16)
where the boundary between the material and void domains is included in the material domain. Using the characteristic function $\chi$, the topology optimization problem can be formulated as follows:
[…]
In topology optimization, the problem described in (13) is reformulated as a material-distribution problem. This allows for topological changes, such as an increase or decrease in the number of holes, during the optimization procedure.
However, topology optimization problems are commonly ill-posed Allaire (2001); therefore, the space of admissible design should incorporate relaxation or regularization techniques to render the problem well-posed. A typical method based on the relaxation of space of an admissible design is the homogenization method Bendsøe and Kikuchi (1988). In this study, we used a level set-based topology optimization method Yamada et al. (2010) to transform an ill-posed problem into a well-posed problem. In this method, the boundary surface of the material domain is represented by the isosurface of a scalar function called the level set function, and changes in the level set function represent changes in the shape of the material domain. The topology optimization problem is regularized by ensuring the proper smoothness of the level set function. This method is described in subsection 2.3.1.
2.3.1 Level set-based topology optimization
In the level set-based method, the scalar function $\phi(\bm{x})\in H^{1}(\Omega_{D})$, also called the level set function, illustrated in the following equation is introduced to represent the shape:
$\left\{\begin{array}[]{lll}-1\leq\phi(\bm{x})<0&\mathrm{for}\quad\bm{x}\in\Omega_{D}\backslash\Omega\\ \phi(\bm{x})=0&\mathrm{for}\quad\bm{x}\in\partial\Omega\\ 1\geq\phi(\bm{x})>0&\mathrm{for}\quad\bm{x}\in\Omega,\\ \end{array}\right.$ (21)
where $\partial\Omega$ denotes the boundaries between the material and void domains. We redefined the characteristic function using the level set function as follows:
$\chi_{\phi}(\phi(\bm{x})):=\left\{\begin{array}[]{ll}1\qquad\mathrm{for}\quad\phi(\bm{x})\geq 0\\ 0\qquad\mathrm{for}\quad\phi(\bm{x})<0.\\ \end{array}\right.$ (24)
In the level set-based method, the topology optimization problem is formulated using the characteristic function defined in (24). However, as previously discussed, optimization that employs characteristic functions as design variables constitutes an ill-posed problem. Consequently, the level set-based method offers regularization to this optimization challenge, which is formulated as follows:
$\begin{split}\underset{\phi}{\text{inf}}\qquad&F_{R}(u,\phi)=\int_{\Omega_{D}}f(u)\chi_{\phi}(\phi)\mathrm{d\Omega}+\int_{\Omega_{D}}\frac{1}{2}\tau|\nabla\phi|^{2}\ \text{d}\Omega\\ \text{subject to}\qquad&\text{the governing equation system},\\ &\text{constraint equations},\end{split}$ (25)
where $\tau\in\mathbb{R}_{+}$ is the regularization coefficient.
Next, we describe a method for updating the level set function. Assuming that the level set function is a function of fictitious time $t$, it is updated by the reaction-diffusion equation as follows:
$\begin{split}\frac{\partial\phi(\bm{x},t)}{\partial t}&=-K(\phi)\left\{-\tilde{c}F^{\prime}-\tau\nabla^{2}\phi(\bm{x},t)\right\},\\ \tilde{c}:&=\frac{c\int_{\Omega_{D}}\ \text{d}\Omega}{\int_{\Omega_{D}}|F^{\prime}|\ \text{d}\Omega},\end{split}$ (26)
where $K\in\mathbb{R}_{+}$ is the proportionality coefficient, $c\in\mathbb{R}_{+}$ is the normalization coefficient, and $F^{\prime}$ is the design sensitivity. In this study, we set $K=1.0$ and $c=2.0$.
3 Formulation of constraints for manufacturability
In this section, we describe the formulation of prerequisites related to manufacturing processes. This study set forth specific manufacturing criteria for piezoelectric energy harvesters, emphasizing configurations amenable to etching methodologies, predominantly utilizing the deep reactive ion etching (DRIE) apparatus Murakami et al. (2013) Laermer (1996). To this end, constraints were introduced to ensure structural uniformity in the thickness direction for both the piezoelectric and substrate domains. Such constraints prevent the emergence of configurations with apertures oriented in oblique or orthogonal directions relative to the processing trajectory. In addition, we advocated the imposition of a substrate-dependent constraint on the piezoelectric domain, precluding the occurrence of piezoelectric materials in areas devoid of a substrate. This constraint ensures that the piezoelectric materials do not exist in areas where the substrate is absent. The intricacies of each constraint are described below.
3.1 Constraint: the same cross-sectional shape in each domain
Within the framework of topology optimization obtained via the level set method, qualitative control over complexity can be regulated by the regularization coefficient $\tau$, as defined in Eq. (25) Yamada et al. (2010). To leverage this control over complexity, we propose strategies to assign regularization coefficients that curtail the complexity in the thickness direction of the structure, thereby facilitating manufacturability via etching. We provided the regularization coefficient as a tensor $\bm{\tau}$, and modified Eq. (26) as follows:
$\begin{split}\frac{\partial\phi(\bm{x},t)}{\partial t}=-K(\phi)\left\{-\tilde{c}F^{\prime}-{\nabla}^{T}\bm{\tau}{\nabla}\phi(\bm{x},t)\right\},\\ \end{split}$ (27)
where
$\bm{\tau}=\left[\begin{array}[]{lll}\tau_{x}&0&0\\ 0&\tau_{y}&0\\ 0&0&\tau_{z}\end{array}\right].$ (31)
In this context, augmenting the component of the regularization coefficient tensor associated with the gradient in a specific direction of the level set function simplifies the complexity along that direction. By defining the thickness direction as $z$ and assigning $\tau_{x},\tau_{y}\ll\tau_{z}$, we can obtain a structure exhibiting minimal complexity in the thickness direction, thus enhancing manufacturability via etching.
Although the method described above controls the complexity of the structure throughout the design domain, the substrate and piezoelectric components can be readily processed via etching, even if their structures differ. It is desirable to control the complexity in the thickness direction within the substrate and piezoelectric domains without affecting the structural differences at the interface of each domain. To this end, in this study, we defined separate level set functions for the substrate and piezoelectric structures, invoking the extended level set method Noda et al. (2022), which is capable of handling multiple materials and updating each level set function concurrently. The fixed design domains for the substrate and piezoelectric material are denoted as $D_{sb}$ and $D_{pe}$, respectively, assuming that these domains remain invariant throughout the optimization procedure. The fixed design domain, denoted as $D$, encompasses both $D_{pe}$ and $D_{sb}$. Within the fixed design domain $D$, two distinct level set functions $\phi_{pe}$ and $\phi_{sb}$ along with characteristic functions $\chi_{ps}$ and $\chi_{sp}$ signifying the respective domains are defined as follows:
$\left\{\begin{array}[]{ll}0<\phi_{pe}(\bm{x})\leq 1&\mathrm{if}~{}~{}\bm{x}\in\Omega_{pe}\\ \phi_{pe}(\bm{x})=0&\mathrm{if}~{}~{}\bm{x}\in\partial\Omega_{pe}\cup D_{sb}\\ -1\leq\phi_{pe}(\bm{x})<0&\mathrm{if}~{}~{}\bm{x}\in D_{pe}\setminus{\Omega}_{pe},\end{array}\right.$ (35)
$\left\{\begin{array}[]{ll}0<\phi_{sb}(\bm{x})\leq 1&\mathrm{if}~{}~{}\bm{x}\in\Omega_{sc}\\ \phi_{sb}(\bm{x})=0&\mathrm{if}~{}~{}\bm{x}\in\partial\Omega_{sc}\cup D_{pe}\\ -1\leq\phi_{sb}(\bm{x})<0&\mathrm{if}~{}~{}\bm{x}\in D_{sb}\setminus{\Omega}_{sb},\end{array}\right.$ (39)
$\left\{\begin{array}[]{ll}\chi_{ps}=1&\mathrm{in}~{}~{}D_{pe}\\ \chi_{ps}=0&\mathrm{in}~{}~{}D_{sb},\end{array}\right.$ (42)
$\left\{\begin{array}[]{ll}\chi_{sp}=0&\mathrm{in}~{}~{}D_{pe}\\ \chi_{sp}=1&\mathrm{in}~{}~{}D_{sb}.\end{array}\right.$ (45)
Each level set function for the piezoelectric and substrate domains is updated as follows:
$\frac{\partial\phi_{pe}}{\partial t}=-K(-\tilde{c}{F}^{\prime}_{pe}-\nabla(\bm{\tau}_{p}\nabla\phi_{pe})),$ (46)
$\frac{\partial\phi_{sb}}{\partial t}=-K(-\tilde{c}{F}^{\prime}_{sb}-\nabla(\bm{\tau}_{sb}\nabla\phi_{sb})),$ (47)
where ${F}^{\prime}_{pe}$ is the gradient of the objective function $F_{pe}$ with respect to $\phi_{pe}$; ${F}^{\prime}_{sb}$ is the gradient of the objective function $F_{sb}$ with respect to $\phi_{sb}$; and $\bm{\tau}_{pe}$ and $\bm{\tau}_{sb}$ are the regularization factor tensors for $\phi_{pe}$ and $\phi_{sb}$, respectively. By individually defining the regularization coefficient tensors for the piezoelectric and substrate domains, the complexity in the thickness direction for each domain can be controlled, thereby enhancing the manufacturability.
3.2 Constraint: no piezoelectric material is placed without a substrate
Herein, we elucidate an approach that allows the imposition of geometric constraints using a fictitious physical function field denoted by $p$. The concept of fictitious physical models described by a steady-state anisotropic advection-diffusion equation was initially introduced by Sato et al. Sato et al. (2017). A further elaboration of the fictitious physical model using the anisotropic diffusion equation was provided by Yamada et al. Yamada and Noguchi (2022), which is described below.
$\begin{split}-\text{div}(\bm{\kappa}_{p}\nabla p)+\chi_{sp}(p-\bar{p}_{s}(2\chi(\phi_{sb})-1))=0&\qquad\text{in}\quad D,\\ p=-\bar{p}_{0}&\qquad\text{on}\quad\Gamma_{p},\\ \nabla p\cdot\bm{n}=0&\qquad\text{on}\quad\partial D\backslash\Gamma_{p},\end{split}$ (48)
$\bm{\kappa}_{p}=\left[\begin{array}[]{lll}\kappa_{x}&0&0\\ 0&\kappa_{y}&0\\ 0&0&\kappa_{z}\end{array}\right],$ (52)
where $\bm{\kappa}_{p}$ represents the conductivity tensor of $p$; $\kappa_{x}$, $\kappa_{y}$, and $\kappa_{z}$ represent the anisotropic conductivities along the $x$, $y$, and $z$ directions, respectively; and $\bar{p}_{s}\in\mathbb{R}_{+}$ and $\bar{p}_{0}\in\mathbb{R}_{+}$ are the constants for the generation and boundary conditions of $p$. In this study, the thickness direction was represented by the $z$-axis, which was employed as a constraint to determine the presence or absence of the substrate. The anisotropic conductivities $\kappa_{x}$ and $\kappa_{y}$ were set to values substantially lower than $\kappa_{z}$. The boundary $\Gamma_{p}$, subjected to the Dirichlet boundary condition, was identified as the boundary opposite to domain $D_{pe}$ within the domain $D_{sb}$. The second term in the initial equation of (48) ensures that $p=\bar{p}_{s}$ within the substrate material domain $\Omega_{sb}$. Within the substrate material domain $\Omega_{sb}$, the fictitious heat $p$ is generated and propagated in the $z$-direction towards the domain $D_{pe}$ when a material is present within $\Omega_{sb}$. Conversely, in the absence of a material in $\Omega_{sb}$, the boundary condition $p=-p_{0}$ is conveyed to the domain $D_{pe}$ owing to the lack of $p$ generation in $\Omega_{sb}$.
3.3 Formulation of the optimal design with proposed constraints
By incorporating the aforementioned methodologies, the optimal design problem is formulated as follows:
$\begin{split}\underset{\phi_{pe}}{\text{inf}}\qquad&F_{pe}(\omega_{oc},\omega_{sc})\\ \underset{\phi_{sb}}{\text{inf}}\qquad&F_{sb}(\omega_{oc},\omega_{sc})\\ \text{subject}\;\text{to}\qquad&a_{oc}(\phi_{pe},\phi_{sb},\omega_{oc},\bm{u}_{oc},\bm{v}_{oc},\varphi_{oc},p)=0,\\ &b_{oc}(\phi_{pe},\phi_{sb},\bm{u}_{oc},\varphi_{oc},v_{\varphi_{oc}},p)=0,\\ &a_{sc}(\phi_{pe},\phi_{sb},\omega_{sc},\bm{u}_{sc},\bm{v}_{sc},p)=0,\\ &a_{p}(\phi_{pe},\phi_{sb},p,v_{p})=0,\\ &G_{V}(\phi_{pe},\varphi_{oc})\leq 0,\end{split}$ (53)
$\begin{split}&a_{oc}(\phi_{pe},\phi_{sb},\omega_{oc},\bm{u}_{oc},\bm{v}_{oc},\varphi_{oc})\\ &\quad=\biggl{(}\int_{\Omega_{pe}}\bm{s}(\bm{v}_{oc})^{T}\bm{C}^{E}_{pe}\bm{s}(\bm{u}_{oc})\ \text{d}\Omega-\omega_{oc}^{2}\int_{\Omega_{pe}}\rho_{pe}\bm{v}_{oc}^{T}\bm{u}_{oc}\ \text{d}\Omega\\ &\qquad+\int_{\Omega_{pe}}\bm{s}(\bm{v}_{oc})^{T}\bm{e}^{T}\bm{\nabla}\varphi_{oc}\ \text{d}\Omega\biggr{)}\\ &\qquad+\biggl{(}\int_{\Omega_{sb}}\bm{s}(\bm{v}_{oc})^{T}\bm{C}^{E}_{sb}\bm{s}(\bm{u}_{oc})\ \text{d}\Omega-\omega_{oc}^{2}\int_{\Omega_{sb}}\rho_{sb}\bm{v}_{oc}^{T}\bm{u}_{oc}\ \text{d}\Omega\biggr{)}=0,\end{split}$ (54)
$\begin{split}&b_{oc}(\phi_{pe},\phi_{sb},\bm{u}_{oc},\varphi_{oc},v_{\varphi_{oc}})\\ &\quad=\int_{\Omega_{pe}}\bm{\nabla}v_{\varphi_{oc}}^{T}\bm{e}\bm{s}(\bm{u}_{oc})\ \text{d}\Omega-\int_{\Omega_{pe}}\bm{\nabla}v_{\varphi_{oc}}^{T}\bm{\varepsilon}^{S}\bm{\nabla}\varphi_{oc}\ \text{d}\Omega=0,\end{split}$ (55)
$\begin{split}&a_{sc}(\phi_{pe},\phi_{sb},\omega_{sc},\bm{u}_{sc},\bm{v}_{sc})\\ &\quad=\biggl{(}\int_{\Omega_{pe}}\bm{s}(\bm{v}_{sc})^{T}\bm{C}^{E}_{pe}\bm{s}(\bm{u}_{sc})\ \text{d}\Omega-\omega_{sc}^{2}\int_{\Omega_{pe}}\rho_{pe}\bm{v}_{sc}^{T}\bm{u}_{sc}\ \text{d}\Omega\biggr{)}\\ &\qquad+\biggl{(}\int_{\Omega_{sb}}\bm{s}(\bm{v}_{sc})^{T}\bm{C}^{E}_{sb}\bm{s}(\bm{u}_{sc})\ \text{d}\Omega-\omega_{sc}^{2}\int_{\Omega_{sb}}\rho_{sb}\bm{v}_{sc}^{T}\bm{u}_{sc}\ \text{d}\Omega\biggr{)}=0,\end{split}$ (56)
$\begin{split}&a_{p}(\phi_{pe},\phi_{sb},p,v_{p})=\int_{D}\nabla v_{p}\bm{\kappa}_{p}\nabla p\ \text{d}\Omega+\int_{D}\chi_{sp}(p-\bar{p}_{s}(2\chi(\phi_{sb})-1))v_{p}\ \text{d}\Omega=0,\end{split}$ (57)
$\begin{split}&G_{V}(\phi_{pe},\varphi)=\int_{\Omega_{pe}}\ \text{d}\Omega-\frac{\varepsilon_{z}L_{z}^{2}\int_{\Omega_{pe}}\bm{n}_{z}\cdot\nabla\varphi\ \text{d}\Omega}{\bar{V}_{minV}}\leq 0,\end{split}$ (58)
where $\bm{C}^{E}_{pe}$ and $\bm{C}^{E}_{sb}$ are the elastic tensors of the piezoelectric and substrate materials, respectively. Similarly, $\rho_{pe}$ and $\rho_{sb}$ represent the densities of the piezoelectric and substrate materials, respectively. Additionally, $v_{p}$ acts as a test function for the fictitious physical function $p$. The potential, symbolized by $\varphi$, is obtained from the displacement $\bm{u}$ derived via the modal superposition method. This is distinct from $\varphi_{oc}$, which is determined based on the eigenvector $\bm{u}_{oc}$. A comprehensive exposition on the derivation of the potential $\varphi$ through the modal superposition method can be found in A.
4 Numerical implementation
4.1 Optimization algorithm
The optimization algorithm is implemented as follows:
  • Step 1:
    The initial level set function is set.
  • Step 2:
    The fictitious physical function field $p$ based on Eq. (57) is solved using the finite element method (FEM).
  • Step 3:
    The eigenvector fields $\bm{u}_{oc}$ and $\bm{u}_{sc}$ and eigenvalues $\omega_{oc}$, $\omega_{sc}$, and $\varphi_{oc}$ derived from Eqs. (54)–(56) are solved using the FEM.
  • Step 4:
    The objective functions $F_{pe}$ and $F_{sb}$ derived from Eq. (9) together with the output voltage $V_{E}$ at a frequency $\bar{\omega_{i}}$ and the constraint function $G_{V}$, which is based on Eq. (58), are evaluated.
  • Step 5:
    If the objective functions converge, the optimization procedure is terminated; otherwise, the sensitivity with respect to the objective functions is computed.
  • Step 6:
    The level set function is updated using the time evolution equation given by Eq. (26), and the optimization procedure returns to step 2.
FreeFEM++ Hecht (2012) was used as the FEM solver. In the following subsections, we explain the approximations of the displacement and electrical potential fields and the sensitivity analysis.
4.2 Approximate solution of the displacement and electrical potential fields based on the Eulerian coordinate system
Within the Eulerian coordinate system, the fixed design domain necessitates the generation of finite elements during each iteration of the optimization process. To mitigate the computational expenditure, an ersatz material approach, as delineated in Allaire et al. (2004), was employed. In particular, the void domain was assumed to be a structural material with a relatively small elastic tensor, and the material properties were assumed to be smoothly distributed in the vicinity of the interface. The governing equations, which dictate the displacement field within the FEM, originally pertain to the material domain $\Omega_{D}$. Utilizing the level set, characteristic, and fictitious physical functions delineated in Section 3, coupled with an approximated Heaviside function $h(\phi)$ to be presented subsequently, we extended these equations to encompass the entire fixed design domain $D$ as follows:
$\begin{split}\chi(\phi)&\approx h(\phi),\\ h(\phi):&=\left\{\begin{array}[]{lll}d&\mathrm{for}\quad\phi<-w\\ \left\{\frac{1}{2}+\frac{\phi}{w}\left[\frac{15}{16}-\frac{\phi^{2}}{w^{2}}(\frac{5}{8}-\frac{3}{16}\frac{\phi^{2}}{w^{2}})\right]\right\}(1-d)+d&\mathrm{for}\quad-w\leq\phi\leq w\\ 1&\mathrm{for}\quad w<\phi\\ \end{array},\right.\end{split}$ (59)
${\bm{C}_{\phi ps}^{E}}:$ $=\bm{C}^{E}_{pe}\chi_{ps}h(\phi_{pe})h(p)+\bm{C}^{E}_{sb}\chi_{sp}h(\phi_{sb}),$ (60)
${\bm{e}}_{\phi ps}:$ $=\bm{e}\chi_{ps}h(\phi_{pe})h(p),$ (61)
${\bm{\varepsilon}}^{S}_{\phi ps}:$ $=\varepsilon_{0}\bm{I}(1-(\chi_{ps}h(\phi_{pe})h(p)+\chi_{sp}h(\phi_{sb})))+\bm{\varepsilon}^{S}\chi_{ps}h(\phi_{pe})h(p),$ (62)
${\rho}_{\phi ps}:$ $=\rho_{pe}\chi_{ps}h(\phi_{pe})h(p)+\rho_{sb}\chi_{sp}h(\phi_{sb}),$ (63)
where $\bm{C}^{E}_{\phi ps}$ is the extended elastic tensor, ${\bm{e}}_{\phi ps}$ is the extended piezoelectric constant tensor, ${\bm{\varepsilon}}^{S}_{\phi ps}$ is the extended permittivity tensor, ${\rho}_{\phi ps}$ is the extended density, $w$ is the transition width of the Heaviside function, and $d$ is a sufficiently small positive number. In this study, we set $w=0.9$ and $d=0.01$.
4.3 Sensitivity analysis
Here, we describe a procedure for determining the sensitivity to update level set functions. In this study, based on Eqs. (54)–(58), we defined Lagrangian $L_{pe}$ and $L_{sb}$ with respect to $\phi_{pe}$ and $\phi_{sb}$, respectively, as follows:
$\begin{split}L_{pe}=&F_{pe}(\omega_{oc},\omega_{sc})+a_{oc}(\phi_{pe},\phi_{sb},\omega_{oc},\bm{u}_{oc},\bm{v}_{ocpe},\varphi_{oc})+b_{oc}(\phi_{pe},\phi_{sb},\bm{u}_{oc},\varphi_{oc},v_{\varphi_{ocpe}})\\ &+a_{sc}(\phi_{pe},\phi_{sb},\omega_{sc},\bm{u}_{sc},\bm{v}_{scpe})+a_{p}(\phi_{pe},\phi_{sb},p,v_{p})+\lambda G_{V}(\phi_{pe},\varphi_{oc}),\end{split}$ (64)
$\begin{split}L_{sb}=&F_{sb}(\omega_{oc},\omega_{sc})+a_{oc}(\phi_{pe},\phi_{sb},\omega_{oc},\bm{u}_{oc},\bm{v}_{ocsb},\varphi_{oc})+b_{oc}(\phi_{pe},\phi_{sb},\bm{u}_{oc},\varphi_{oc},v_{\varphi_{ocsb}})\\ &+a_{sc}(\phi_{pe},\phi_{sb},\omega_{sc},\bm{u}_{sc},\bm{v}_{scsb})+a_{p}(\phi_{pe},\phi_{sb},p,v_{p}),\end{split}$ (65)
where $\bm{v}_{ocpe}$, $\bm{v}_{ocsb}$, $\bm{v}_{sc}$, $\bm{v}_{scpe}$, $\bm{v}_{scsb}$, $v_{\varphi_{ocpe}}$, $v_{\varphi_{ocsb}}$, and $\lambda$ are the Lagrange multipliers. Using the adjoint method, sensitivity was determined as follows:
$\begin{split}{F}^{\prime}_{pe}=&c_{ocpe}\left\{a_{oc}(1,\phi_{sb},\omega_{oc},\bm{u}_{oc},\bm{u}_{oc},\varphi_{oc})+b_{oc}(1,\bm{u}_{oc},\varphi_{oc},\varphi_{oc})\right\}\\ &+c_{scpe}a_{sc}(1,\phi_{sb},\omega_{sc},\bm{u}_{sc},\bm{u}_{sc})+\lambda,\end{split}$ (66)
$\begin{split}{F}^{\prime}_{sb}=&c_{ocsb}\left\{a_{oc}(\phi_{pe},1,\omega_{oc},\bm{u}_{oc},\bm{u}_{oc},\varphi_{oc})\right\}\\ &+c_{scsb}\left\{a_{sc}(\phi_{pe},1,\omega_{sc},\bm{u}_{sc},\bm{u}_{sc})\right\}.\end{split}$ (67)
The sensitivity analysis is explained in detail in B.
5 Numerical examples
In this section, several numerical examples are presented to demonstrate the utility and validity of the proposed method.
The design domain and associated boundary conditions are illustrated in Figure 2. Let $D_{pe}$ denote a square, colored in blue in the figure, with a side length of 500 mm and a thickness of 4 mm. The object in this domain was a piezoelectric material. Similarly, let $D_{sb}$ symbolize a square, indicated in grey in the figure, with a side length of 500 mm and a thickness of 36 mm. The object within this domain did not exhibit any piezoelectric properties. The domain measuring 500 mm in length, 20 mm in width, and 36 mm in thickness, depicted in dark purple in the figure, was a nondesigned domain inhabited by a nonpiezoelectric material. Correspondingly, the domain that is 500 mm in length, 20 mm in breadth, and 4 mm in thickness, illustrated in light purple in the figure, represents a nondesigned domain occupied by a piezoelectric material. A fixed boundary condition was imposed on the surfaces adjacent to as well as the opposite domain $D_{sb}$. The domain colored red on the right side of the figure, delineating a 20-mm square with a thickness of 40 mm, is designated as the weight domain, which is a nondesigned domain filled with objects devoid of piezoelectric properties. The density was assumed to be 100 times that of the substrate material within this weight domain. The number of modes to be contemplated was fixed at $n=4$, and the specified natural frequencies were $[\bar{\omega}_{0},\bar{\omega}_{1},\bar{\omega}_{2},\bar{\omega}_{3}]=[70\;\text{Hz},435\;\text{Hz},450\;\text{Hz},500\;\text{Hz}]$. The initial structure was conjectured to be such that $\phi_{pe}=1$ in $D_{pe}$ and $\phi_{sb}=1$ in $D_{sb}$, indicating that the entire fixed design domain was occupied by the material domain.
Figure 2: Design domain and boundary conditions.
5.1 Effectiveness of the proposed constraints for manufacturability
We conducted numerical experiments under nine conditions, as listed in Table 1. Each condition was characterized by a distinct setting of the regularization factor $\tau_{z}$ in the thickness direction, the representation of level set functions (either a singular function $\phi$ or a pair of functions $\phi_{pe}$ and $\phi_{sb}$), and the status of constraint implementation via a fictitious field $p$. Specifically, for each condition, the regularization factor $\tau_{z}$ was set to 1.0×$\times 10^{-6}$, 1.0×$\times 10^{-4}$, or 1.0×$\times 10^{-2}$. Furthermore, the level set functions were either represented solely by $\phi$ or by a combination of $\phi_{pe}$ and $\phi_{sb}$, and the constraints were applied using the fictitious field $p$ for some conditions. Each of these nine configurations was evaluated through 1000 iterations of computation. The parameters corresponding to each condition, coupled with the average values of the objective function $F_{k}$ and $F_{\omega}$ taken over the last 100 iterations, are also presented in Table 1. Under all conditions, $\alpha_{pe}$ and $\alpha_{sb}$ were set to $0.95$. The resulting structures for each condition are shown in Figure 3.
The tradeoff relationship between the objective functions $F_{k}$ and $F_{\omega}$ for conditions (a)–(i) delineated in Table 1 is depicted in Figure 4. From the figure, a clear tradeoff relationship can be discerned between $F_{\omega}$ and $F_{k}$, the balance of which is influenced by changes in the parameter $\tau_{z}$. Moreover, conditions (d)–(f) yielded smaller $F_{k}$ values than conditions (a)–(c), given the same $\tau_{z}$ value. This phenomenon is attributed to the fact that, in conditions (d)–(f), the architecture of the piezoelectric film remained uninfluenced by the structure of the substrate, and, apart from regularization with respect to $\phi_{pe}$, it does not adhere to any constraints. This allows for a wider variety of potential solutions, thereby enhancing the structural degrees of freedom of piezoelectric films. In conditions (g)–(i), although the structure of the piezoelectric film was influenced by the structure of the substrate, the converse does not hold true, resulting in $F_{\omega}$ values that are lower than those under conditions (d)–(f).
Table 1: Summary of parameter settings and objective functions: regularization factor $\tau_{z}$ in the thickness direction, the representation of level set functions, and the status of constraint implementation via a fictitious field $p$ and the corresponding objective functions $F_{k}$ and $F_{\omega}$ for each condition.
Objective function
condition $\tau_{z}$ level set function $p$ constraint $F_{k}$ $F_{\omega}$
(a) 1.0×$\times 10^{-6}$ $\phi$ not constrained 4.47×$\times 10^{-5}$ 1.52×$\times 10^{-5}$
(b) 1.0×$\times 10^{-4}$ $\phi$ not constrained 7.07×$\times 10^{-5}$ 6.70×$\times 10^{-6}$
(c) 1.0×$\times 10^{-2}$ $\phi$ not constrained 8.43×$\times 10^{-5}$ 4.98×$\times 10^{-6}$
(d) 1.0×$\times 10^{-6}$ $\phi_{pe}\;$ and $\;\phi_{sb}$ not constrained 4.50×$\times 10^{-5}$ 1.51×$\times 10^{-5}$
(e) 1.0×$\times 10^{-4}$ $\phi_{pe}\;$ and $\;\phi_{sb}$ not constrained 5.80×$\times 10^{-5}$ 1.33×$\times 10^{-5}$
(f) 1.0×$\times 10^{-2}$ $\phi_{pe}\;$ and $\;\phi_{sb}$ not constrained 7.20×$\times 10^{-5}$ 1.08×$\times 10^{-5}$
(g) 1.0×$\times 10^{-6}$ $\phi_{pe}\;$ and $\;\phi_{sb}$ $p$ constrained 4.70×$\times 10^{-5}$ 1.25×$\times 10^{-5}$
(h) 1.0×$\times 10^{-4}$ $\phi_{pe}\;$ and $\;\phi_{sb}$ $p$ constrained 4.53×$\times 10^{-5}$ 1.28×$\times 10^{-5}$
(i) 1.0×$\times 10^{-2}$ $\phi_{pe}\;$ and $\;\phi_{sb}$ $p$ constrained 8.14×$\times 10^{-5}$ 4.79×$\times 10^{-6}$
Figure 3: Optimal configuration of an unimorph cantilevered energy harvester for conditions (a)–(i) listed in Table 1. The red region denotes the piezoelectric material domain $\Omega_{pe}$, and the blue region represents the substrate material domain $\Omega_{sb}$.
Figure 4: Tradeoff relationship between the objective functions $F_{k}$ and $F_{\omega}$ under conditions (a)–(i) listed in Table 1.
For each condition, we computed two normalized quantities: $N_{\phi 1}$ and $N_{\phi 2}$. $N_{\phi 1}$ is the proportion of nodes at which the product of the level set functions with their adjacent nodes in the z-negative direction is negative, suggesting a boundary with the adjacent nodes. Conversely, $N_{\phi 2}$ is the proportion of nodes not located at the boundaries of domains $D_{sb}$ and $D_{pe}$. These values were normalized by the total number of nodes. The computed values for $N_{\phi 1}$ and $N_{\phi 2}$ for each condition are provided in Table 2.
As delineated in Table 2, irrespective of the level set function configuration and the existence of a $p$ constraint, an increase in the $\tau_{z}$ value corresponded to a reduction in the z-directional complexity. When two level set functions were utilized, as in conditions (d)–(i), the $N_{\phi 1}$ value surpassed that in conditions (a)–(c), where a single level set function embodied the entirety of the design domain. Contrastingly, $N_{\phi 2}$, excluding the interface between $D_{sb}$ and $D_{pe}$, did not exhibit a significant difference from $N_{\phi 1}$ under conditions (a)–(c) but was less than $N_{\phi 1}$ in conditions (d)–(i). This phenomenon can be attributed to the fact that, when two level set functions are employed, the structural complexity is not restrained at the boundary between domains $D_{sb}$ and $D_{pe}$, even upon altering the $\tau_{z}$ value. Hence, the $N_{\phi 1}$ value was considered higher than that in conditions (a)–(c).
Table 2: Number of nodes where the product of their level set function values with nodes adjacent to the negative direction of the z-axis is negative for each condition
condition $\tau_{z}$ $N_{\phi 1}$ $N_{\phi 2}$
(a) 1.0×$\times 10^{-6}$ 0.08818 0.07629
(b) 1.0×$\times 10^{-4}$ 0.01965 0.01819
(c) 1.0×$\times 10^{-2}$ 0.00049 0.00042
(d) 1.0×$\times 10^{-6}$ 0.08890 0.07138
(e) 1.0×$\times 10^{-4}$ 0.05058 0.03523
(f) 1.0×$\times 10^{-2}$ 0.01898 0.00079
(g) 1.0×$\times 10^{-6}$ 0.14124 0.10782
(h) 1.0×$\times 10^{-4}$ 0.08660 0.06932
(i) 1.0×$\times 10^{-2}$ 0.00340 0.00200
5.2 Optimal configuration subjected to a minimum output voltage constraint
Herein, the efficacy of the introduced constraint with respect to output voltage is critically examined. Consistently across all scenarios, the substrate-dependent constraint was enforced utilizing the fictitious physical function $p$. Moreover, the material domain was characterized by dual-level set functions, $\phi_{pe}$ and $\phi_{sb}$, with the parameter $\tau_{z}$ assigned a value of $1.0\times 10^{-2}$. Table 3 enumerates the parameters for the minimum output voltage constraint under various conditions, alongside their respective objective functions $F_{k}$, $F_{\omega}$, and output voltage $V_{E}$. The optimal configurations for each scenario are graphically represented in Figure 5.
Table 3: Summary of constraint parameters and objective functions: minimum output voltage constraint $\bar{V}_{minV}$ and the corresponding objective functions $F_{k}$ and $F_{\omega}$, the output voltage $V_{E}$, and the volume of piezoelectric material domain $\Omega_{pe}$ for each condition.
Constraint Objective function Output voltage Volume of $\Omega_{pe}$
condition $\bar{V}_{minV}$ $F_{k}$ $F_{\omega}$ $V_{E}$ $V_{pe}$
(j) 9.0×$\times 10^{-3}$ 8.44×$\times 10^{-5}$ 4.56×$\times 10^{-6}$ 9.73×$\times 10^{-3}$ 4.47 ×$\times 10^{5}$
(k) 9.5×$\times 10^{-3}$ 1.82×$\times 10^{-4}$ 4.87×$\times 10^{-6}$ 9.50×$\times 10^{-3}$ 2.31 ×$\times 10^{5}$
(l) 1.0×$\times 10^{-2}$ 2.03×$\times 10^{-4}$ 5.60×$\times 10^{-6}$ 1.00×$\times 10^{-2}$ 2.07 ×$\times 10^{5}$
(m) 1.05×$\times 10^{-2}$ 2.21×$\times 10^{-4}$ 6.23×$\times 10^{-6}$ 1.05×$\times 10^{-2}$ 1.90 ×$\times 10^{5}$
(n) 1.1×$\times 10^{-2}$ 2.34×$\times 10^{-4}$ 6.72×$\times 10^{-6}$ 1.10×$\times 10^{-2}$ 1.80 ×$\times 10^{5}$
(o) 1.15×$\times 10^{-2}$ 2.45×$\times 10^{-4}$ 7.11×$\times 10^{-6}$ 1.15×$\times 10^{-2}$ 1.71 ×$\times 10^{5}$
(p) 1.2×$\times 10^{-2}$ 2.39×$\times 10^{-4}$ 1.02×$\times 10^{-5}$ $\times 10^{-2}$ 1.66 ×$\times 10^{5}$
(q) 1.3×$\times 10^{-2}$ 2.61×$\times 10^{-4}$ 1.11×$\times 10^{-5}$ 1.30×$\times 10^{-2}$ 1.49 ×$\times 10^{5}$
Figure 5: Optimal configuration of the unimorph cantilevered energy harvester for conditions (j)–(q), as detailed in Table 3. The red region denotes the piezoelectric material domain $\Omega_{pe}$, and the blue region represents the substrate material domain $\Omega_{sb}$.
As shown in Table 3, the output voltage $V_{E}$ obtained under condition (j) exceeded the stipulated minimum voltage constraint. However, under the other conditions, the output voltages were aligned exactly with their respective constraints highlighted the effectiveness and precision of the minimum output voltage constraint. Furthermore, as the output voltage constraint values increased, there was a concomitant reduction in the volume of the piezoelectric material domain, represented by $\Omega_{pe}$. As elucidated by Eq. (12), this reduction in the volume of the piezoelectric material domain results in a diminished capacitance, which, in turn, amplifies the production voltage. Concomitantly, an increase in the $\bar{V}_{minV}$ value invariably leads to a corresponding increase in both objective functions $F_{k}$ and $F_{\omega}$. It is cogent to infer that imposing a minimum output voltage constraint invariably curtails the potential for optimizing the objective functions $F_{k}$ and $F_{\omega}$.
As shown in Figure 5, under conditions (k)–(q) with an active minimum output voltage constraint, the piezoelectric material domain was predominantly located near the fixed boundary and at the midpoint of the beam. Proximate to the fixed boundary of the beam, the aggregation of the piezoelectric material domain can be attributed to the stress concentration observed in the first-order eigenmode. The region encompassing the piezoelectric material domain located around the center of the beam was aligned with the zone of maximal displacement in the second-order eigenmodes.
6 Conclusion
In this study, we proposed a method for designing piezoelectric energy harvesters to improve their manufacturability. The proposed method derives both the substrate and piezoelectric designs that maximize the electromechanical coupling coefficient. This method allows the eigenfrequency of the device and minimum output voltage to be set to the desired values. Furthermore, the proposed method yields a design that can be manufactured using a microfabrication process. The results of this study are summarized as follows:
  • 1.
    A topology optimization method was formulated to achieve the desired values for the eigenfrequency and minimum output voltage while concurrently maximizing the electromechanical coupling coefficient. An optimization algorithm was constructed for the numerical analysis.
  • 2.
    Cross-sectional shape constraints and substrate-dependent constraints were formulated to improve manufacturability with microfabrication.
  • 3.
    Several numerical examples were provided to demonstrate the utility and validity of the proposed method. The proposed method can provide solutions that satisfy the cross-sectional shape, substrate-depend, and minimum output voltage constraints.
6.1 Limitations and future work
Despite the promising results, this study has some limitations. As the constraints became more stringent, the value of the objective function deteriorated. This indicates that the proposed method may struggle to provide solutions with significant electrical-mechanical coupling coefficients under highly restrictive conditions.
In future work, we aim to actually manufacture devices based on the designs obtained through the proposed method. This will allow us to validate the practicality of the solutions obtained and further demonstrate the utility of the proposed method.
6.2 Final remarks
The proposed method for designing piezoelectric energy harvesters has shown significant potential for improving their manufacturability. We believe that this research contributes to the field of energy harvesting and opens up new possibilities for the design of efficient and manufacturable energy harvesters.
Appendix A Computation of the output voltage using the modal superposition method
In this study, output voltage was computed using the modal superposition method. In A.1, we explain the method used to solve the eigenvalue problem using the finite element method(FEM). In A.2, we explain the method used to yield the output voltage using the modal superposition method based on the results obtained in A.1.
A.1 Eigenvalue and eigenvector determination
In this subsection, we elucidate an approach to solve the eigenvalue problem for piezoelectric materials using the FEM. Under open-circuit conditions, the relationships can be expressed in the form of finite element equations, as described below.
$(\bm{K}+j\omega\bm{R}-\omega_{oc}^{2}\bm{M})\bm{u}+\bm{P}\varphi$ $=0,$ (68)
$\bm{P}^{T}\bm{u}-\bm{G}\varphi$ $=0,$ (69)
wherein $j$ denotes the imaginary unit; $\bm{u}$ and $\varphi$ correspond to the displacement and electric potential under open-circuit conditions, respectively; and $\bm{P}$ and $\bm{G}$ denote the piezoelectric and dielectric matrices, respectively. Notably, $\bm{R}$ signifies the damping coefficient matrix inherent to the system. In the absence of an external charge, denoted by $\bm{q}$, and an external force, symbolized by $\bm{f}$, the potential $\varphi$ can be eliminated. This leads to:
$\begin{split}&\bm{P}^{T}\bm{u}=\bm{G}\varphi,\\ &\varphi=\bm{G}^{-1}\bm{P}^{T}\bm{u},\end{split}$ (70)
$(\bm{K}+j\omega\bm{R}-\omega^{2}\bm{M}+\bm{PG}^{-1}\bm{P}^{T})\bm{u}=0.$ (71)
Consequently, the eigenvalue equation pertinent to the open-circuit conditions can be expressed as follows:
$\bm{M}^{-1}(\bm{K}+\bm{PG}^{-1}\bm{P}^{T})\bm{u}_{oc}=\omega_{oc}^{2}\bm{u}_{oc},$ (72)
where $\bm{u}_{oc}$ represents the eigenvector associated with the eigenvalue $\omega_{oc}$.
Conversely, for a short-circuited electrode, the potential difference between the electrodes is zero. By postulating that both the external charge $\bm{q}$ and the external force $\bm{f}$ are zero, analogous to the open-circuit condition, and by grounding the electrode to establish a zero potential, Eq. (6) yields the following equation:
$(\bm{K}+j\omega\bm{R}-\omega^{2}\bm{M})\bm{u}=0.$ (73)
The eigenvalue equation for the short-circuit condition is given as follows:
$\bm{M}^{-1}\bm{K}\bm{u}_{sc}=\omega_{sc}^{2}\bm{u}_{sc},$ (74)
where $\bm{u}_{sc}$ represents the eigenvector corresponding to the eigenvalue $\omega_{sc}$.
A.2 Evaluation of displacement and output voltage using the modal superposition method
Because the eigenvectors were normalized, it was necessary to consider the magnitude of displacement under forced vibration conditions to evaluate the output voltage. To achieve this, we utilized the modal superposition method to determine the displacement. First, the equation of motion is expressed as,
$\bm{M}\ddot{\bm{u}}+\bm{C}\dot{\bm{u}}+\bm{Ku}=\bm{f}.$ (75)
For undamped free vibrations, we determined the eigenfrequency $\omega_{ocr}$ and the eigenvector $\bm{u}_{oci}$. Using the modal transformation matrix $\bm{\Phi}$, we performed a coordinate transformation such that $\bm{u}=\bm{\Phi}\bm{q}$. The modal transformation matrix $\bm{\Phi}$ is given by
$\bm{\Phi}=\left[\bm{u}_{oc1},\bm{u}_{oc2},\bm{u}_{oc3},\cdots,\bm{u}_{ocn}\right].$ (76)
Multiplying Eq. (75) on the left with $\bm{\Phi}^{T}$, we obtain:
$\bm{\Phi}^{T}\bm{M}\bm{\Phi}\ddot{\bm{q}}+\bm{\Phi}^{T}\bm{C}\bm{\Phi}\dot{\bm{q}}+\bm{\Phi}^{T}\bm{K}\bm{\Phi}\bm{q}=\bm{\Phi}^{T}\bm{f}.$ (77)
Through normalization, we transform each coefficient term as,
$\bm{\Phi}^{T}\bm{M}\bm{\Phi}$ $=\bm{I},$ (78)
$\bm{\Phi}^{T}\bm{C}\bm{\Phi}=\bm{D}$ $=\left[\begin{array}[]{llll}2\zeta_{1}\omega_{1}&0&\cdots&0\\ 0&2\zeta_{2}\omega_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&2\zeta_{n}\omega_{n}\end{array}\right],$ (83)
$\bm{\Phi}^{T}\bm{K}\bm{\Phi}$ $=\left[\begin{array}[]{llll}\omega_{1}^{2}&0&\cdots&0\\ 0&\omega_{2}^{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\omega_{n}^{2}\end{array}\right],$ (88)
$\bm{\Phi}^{T}\bm{f}$ $=\left[\begin{array}[]{llll}F_{1}(t)&0&\cdots&0\\ 0&F_{2}(t)&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&F_{n}(t)\end{array}\right].$ (93)
Assuming that the damping is small and the off-diagonal terms of the damping matrix are negligible compared with the diagonal terms, the transformation allowed us to express the motion Eq. (75) as $n$ independent equations for each eigenmode as,
$\ddot{q}_{i}(t)+2\zeta_{i}\omega_{i}\dot{q}_{i}+\omega_{i}^{2}q_{i}(t)=F_{i}(t)\qquad(i=1,2,\cdots,n),$ (94)
$q_{i}(t)=\frac{F_{i}(t)}{\omega_{i}\sqrt{\left(1-(\bar{\omega}_{i}/\omega_{i})^{2}\right)^{2}+4\zeta_{i}^{2}(\bar{\omega}_{i}/\omega_{i})^{2}}}.$ (95)
From the above , displacement is yielded as follows:
$\bm{u}(t)=\sum_{i=1}^{n}q_{i}(t)\bm{u}_{oci}.$ (96)
Incorporating this $\bm{u}$ into Eq. (70) yields the potential $\varphi$. Notably, $\zeta_{i}=0.01$ for $i=1,2,\cdots n$. Finally, incorporating $\varphi$ into Eq. (10) yields the output voltage.
Appendix B Sensitivity analysis
In this section, we provide a detailed description of the sensitivity analysis using mathematical formulations.
At the stationary point of the Lagrangian, the following optimality conditions hold:
$\cfrac{\partial L_{pe}}{\partial\hat{\bm{u}}_{oc}},\delta\bm{u}_{oc}\right>\middle|_{opt}\right.=\left.\left<\cfrac{\partial(F_{pe}+a_{oc}+b_{oc})}{\partial\hat{\bm{u}}_{oc}},\delta\bm{u}_{oc}\right>\middle|_{opt}\right.=0,$ (97)
$\cfrac{\partial L_{pe}}{\partial\bm{v}_{ocpe}},\delta\bm{v}_{ocpe}\right>\middle|_{opt}\right.=\left.\left<\cfrac{\partial(F_{pe}+a_{oc})}{\partial\bm{v}_{ocpe}},\delta\bm{v}_{ocpe}\right>\middle|_{opt}\right.=0,$ (98)
$\cfrac{\partial L_{pe}}{\partial\hat{\varphi}_{oc}},\delta\varphi_{oc}\right>\middle|_{opt}\right.=\left.\left<\cfrac{\partial(F_{pe}+a_{oc}+b_{oc}+\lambda d_{o})}{\partial\hat{\varphi}_{oc}},\delta\varphi_{oc}\right>\middle|_{opt}\right.=0,$ (99)
$\cfrac{\partial L_{pe}}{\partial v_{\varphi_{ocpe}}},\delta v_{\varphi_{ocpe}}\right>\middle|_{opt}\right.=\left.\left<\cfrac{\partial(F_{pe}+b_{oc})}{\partial v_{\varphi_{ocpe}}},\delta v_{\varphi_{ocpe}}\right>\middle|_{opt}\right.=0,$ (100)
$\left.\cfrac{\partial L_{pe}}{\partial\hat{\omega}_{oci}}\middle|_{opt}\right.=\left.\left<\cfrac{\partial(F_{pe}+a_{oc}(\omega_{oci}))}{\partial\hat{\omega}_{oci}},\delta\omega_{oci}\right>\middle|_{opt}\right.=0,$ (101)
$\cfrac{\partial L_{pe}}{\partial\hat{\bm{u}}_{sc}},\delta\bm{u}_{sc}\right>\middle|_{opt}\right.=\left.\left<\cfrac{\partial(F_{pe}+a_{sc})}{\partial\hat{\bm{u}}_{sc}},\delta\bm{u}_{sc}\right>\middle|_{opt}\right.=0,$ (102)
$\cfrac{\partial L_{pe}}{\partial\bm{v}_{scpe}},\delta\bm{v}_{scpe}\right>\middle|_{opt}\right.=\left.\left<\cfrac{\partial(F_{pe}+a_{sc})}{\partial\bm{v}_{scpe}},\delta\bm{v}_{scpe}\right>\middle|_{opt}\right.=0,$ (103)
$\left.\cfrac{\partial L_{pe}}{\partial\hat{\omega}_{sci}}\middle|_{opt}\right.=\left.\left<\cfrac{\partial(F_{pe}+a_{sc}(\omega_{sci}))}{\partial\hat{\omega}_{sci}},\delta\omega_{sci}\right>\middle|_{opt}\right.=0,$ (104)
$\cfrac{\partial L_{sb}}{\partial\hat{\bm{u}}_{oc}},\delta\bm{u}_{oc}\right>\middle|_{opt}\right.=\left.\left<\cfrac{\partial(F_{sb}+a_{oc}+b_{oc})}{\partial\hat{\bm{u}}_{oc}},\delta\bm{u}_{oc}\right>\middle|_{opt}\right.=0,$ (105)
$\cfrac{\partial L_{sb}}{\partial\bm{v}_{ocsb}},\delta\bm{v}_{ocsb}\right>\middle|_{opt}\right.=\left.\left<\cfrac{\partial(F_{sb}+a_{oc})}{\partial\bm{v}_{ocsb}},\delta\bm{v}_{ocsb}\right>\middle|_{opt}\right.=0,$ (106)
$\cfrac{\partial L_{sb}}{\partial\hat{\varphi}_{oc}},\delta\varphi_{oc}\right>\middle|_{opt}\right.=\left.\left<\cfrac{\partial(F_{sb}+a_{oc}+b_{oc}+\lambda d_{o})}{\partial\hat{\varphi}_{oc}},\delta\varphi_{oc}\right>\middle|_{opt}\right.=0,$ (107)
$\cfrac{\partial L_{sb}}{\partial v_{\varphi_{ocsb}}},\delta v_{\varphi_{ocsb}}\right>\middle|_{opt}\right.=\left.\left<\cfrac{\partial(F_{sb}+b_{oc})}{\partial v_{\varphi{ocsb}}},\delta v_{\varphi{ocsb}}\right>\middle|_{opt}\right.=0,$ (108)
$\left.\cfrac{\partial L_{sb}}{\partial\hat{\omega}_{oci}}\middle|_{opt}\right.=\left.\left<\cfrac{\partial(F_{sb}+a_{oc}(\omega_{oci}))}{\partial\hat{\omega}_{oci}},\delta\omega_{oci}\right>\middle|_{opt}\right.=0,$ (109)
$\cfrac{\partial L_{sb}}{\partial\hat{\bm{u}}_{sc}},\delta\bm{u}_{sc}\right>\middle|_{opt}\right.=\left.\left<\cfrac{\partial(F_{sb}+a_{sc})}{\partial\hat{\bm{u}}_{sc}},\delta\bm{u}_{sc}\right>\middle|_{opt}\right.=0,$ (110)
$\cfrac{\partial L_{sb}}{\partial\bm{v}_{scsb}},\delta\bm{v}_{scsb}\right>\middle|_{opt}\right.=\left.\left<\cfrac{\partial(F_{sb}+a_{sc})}{\partial\bm{v}_{scsb}},\delta\bm{v}_{scsb}\right>\middle|_{opt}\right.=0,$ (111)
$\left.\cfrac{\partial L_{sb}}{\partial\hat{\omega}_{sci}}\middle|_{opt}\right.=\left.\left<\cfrac{\partial(F_{sb}+a_{sc}(\omega_{sci}))}{\partial\hat{\omega}_{sci}},\delta\omega_{sci}\right>\middle|_{opt}\right.=0,$ (112)
where the expressions within the brackets represent the directional derivatives of the functional. The optimality conditions, defined by Eqs. (97)– (112), revealed that the variables $\hat{\bm{u}}_{oc}$, $\hat{\bm{u}}_{sc}$, $\hat{\varphi}_{oc}$, $\hat{\omega}_{oc}$, and $\hat{\omega}_{sc}$ coincides with state the variables $\bm{u}_{oc}$, $\bm{u}_{sc}$, $\varphi_{oc}$, $\omega_{oc}$, and $\omega_{sc}$, respectively.
Eqs. (97), (99), and (101) can be used to derive $\bm{v}_{ocpe}$, $\bm{v}_{ocsb}$, $\bm{v}_{scpe}$, $\bm{v}_{scsb}$, $v_{\varphi{ocpe}}$, and $v_{\varphi{ocsb}}$ as follows:
$\bm{v}_{ocpe}$ $=c_{ocpe}\bm{u}_{oc},\quad\bm{v}_{scpe}=c_{scpe}\bm{u}_{sc},\quad v_{\varphi_{ocpe}}=c_{ocpe}\varphi_{oc},$ (113)
$c_{ocpe}$ $=\frac{-\alpha_{pe}\omega_{sc}^{2}}{(\omega_{oc}^{2}-\omega_{sc}^{2})^{2}}+(1-\alpha_{pe})\sum^{n}_{i=1}\frac{(\omega_{oci}-\bar{\omega}_{i})}{\omega_{oci}\bar{\omega}_{i}^{2}},$ (114)
$c_{scpe}$ $=\frac{\alpha_{pe}\omega_{oc}^{2}}{(\omega_{oc}^{2}-\omega_{sc}^{2})^{2}}.$ (115)
$\bm{v}_{ocsb}$ $=c_{ocsb}\bm{u}_{oc},\quad\bm{v}_{scsb}=c_{scsb}\bm{u}_{sc},\quad v_{\varphi_{ocsb}}=c_{ocsb}\varphi_{oc},$ (116)
$c_{ocsb}$ $=\frac{-\alpha_{sb}\omega_{sc}^{2}}{(\omega_{oc}^{2}-\omega_{sc}^{2})^{2}}+(1-\alpha_{sb})\sum^{n}_{i=1}\frac{(\omega_{oci}-\bar{\omega}_{i})}{\omega_{oci}\bar{\omega}_{i}^{2}},$ (117)
$c_{scsb}$ $=\frac{\alpha_{sb}\omega_{oc}^{2}}{(\omega_{oc}^{2}-\omega_{sc}^{2})^{2}}.$ (118)
From the above equations, the sensitivities for updating $\phi_{pe}$ and $\phi_{sb}$ can be obtained as follows:
$\begin{split}{F}^{\prime}_{pe}=\frac{\partial L_{pe}}{\partial\phi_{pe}}=&c_{ocpe}\left\{\frac{\partial a_{oc}(\phi_{pe},\phi_{sb},\omega_{oc},\bm{u}_{oc},\bm{u}_{oc},\varphi_{oc})}{\partial\phi_{pe}}+\frac{\partial b_{oc}(\phi_{pe},\bm{u}_{oc},\varphi_{oc},\varphi_{oc})}{\partial\phi_{pe}}\right\}\\ &+c_{scpe}\left\{\frac{\partial a_{sc}(\phi_{pe},\phi_{sb},\omega_{sc},\bm{u}_{sc},\bm{u}_{sc})}{\partial\phi_{pe}}\right\}+\lambda,\\ =&c_{ocpe}\left\{a_{oc}(1,\phi_{sb},\omega_{oc},\bm{u}_{oc},\bm{u}_{oc},\varphi_{oc})+b_{oc}(1,\bm{u}_{oc},\varphi_{oc},\varphi_{oc})\right\}\\ &+c_{scpe}a_{sc}(1,\phi_{sb},\omega_{sc},\bm{u}_{sc},\bm{u}_{sc})+\lambda,\end{split}$ (119)
$\begin{split}{F}^{\prime}_{sb}=\frac{\partial L_{sb}}{\partial\phi_{sb}}=&c_{ocsb}\left\{\frac{\partial a_{oc}(\phi_{pe},\phi_{sb},\omega_{oc},\bm{u}_{oc},\bm{u}_{oc},\varphi_{oc})}{\partial\phi_{sb}}+\frac{\partial b_{oc}(\phi_{pe},\bm{u}_{oc},\varphi_{oc},\varphi_{oc})}{\partial\phi_{sb}}\right\}\\ &+c_{scsb}\left\{\frac{\partial a_{sc}(\phi_{pe},\phi_{sb},\omega_{sc},\bm{u}_{sc},\bm{u}_{sc})}{\partial\phi_{sb}}\right\},\\ =&c_{ocsb}\left\{a_{oc}(\phi_{pe},1,\omega_{oc},\bm{u}_{oc},\bm{u}_{oc},\varphi_{oc})\right\}\\ &+c_{scsb}\left\{a_{sc}(\phi_{pe},1,\omega_{sc},\bm{u}_{sc},\bm{u}_{sc})\right\}.\end{split}$ (120)